### v2

# The fractional Helly number for separable convexity spaces

Abstract.
A convex lattice set in $\mathbb{Z}^{d}$ is the intersection of a convex set in $\mathbb{R}^{d}$ and the integer lattice $\mathbb{Z}^{d}$. A well-known theorem of Doignon states that the Helly number of $d$-dimensional convex lattice sets equals $2^{d}$, while a remarkable theorem of Bárány and Matoušek states that the fractional Helly number is only $d+1$. In this paper we generalize their result to abstract convexity spaces that are equipped with a suitable separation property. We also disprove a conjecture of Bárány and Kalai about an existence of fractional Helly property for a family of solutions of bounded-degree polynomial inequalities.
1. Introduction
1.1. The fractional Helly theorem
The well-known theorem of Helly [11] asserts that if $F$ is a finite family of at least $d+1$ convex sets in $\mathbb{R}^{d}$ in which every $(d+1)$-tuple is intersecting, then the entire family $F$ is intersecting. Here, we say that a family of sets is intersecting when the intersection of all its members is nonempty. Among the numerous generalizations and variations of Helly’s theorem, a particularly important one is the fractional Helly theorem, due to Katchalski and Liu [14]. Roughly speaking, it says that if the family contains many intersecting $(d+1)$-tuples, then the family contains a large intersecting subfamily. More precisely, the theorem asserts that if $F$ is a finite family of convex sets in $\mathbb{R}^{d}$ with at least $\alpha\cdot\binom{|F|}{d+1}$ intersecting $(d+1)$-tuples, then $F$ contains an intersecting subfamily of size at least $\beta\cdot|F|$, where $\beta$ is some positive constant which depends only on $d$ and $\alpha$. In fact, the works of Eckhoff [8] and Kalai [13] imply that one may take $\beta=1-(1-\alpha)^{1/(d+1)}$, which also turns out to be optimal.
The importance of the fractional Helly theorem in modern combinatorial convexity was exhibited by Alon and Kleitman in their celebrated proof of the $(p,q)$ theorem [2], and further investigated in e.g. [1]. It turns out that in these applications, it is the existence of $\beta$ that plays the most crucial role, more so than its precise value in terms of $d$ and $\alpha$. See also [17, Chapters 8–10] and [4] for further discussion and references.
1.2. The fractional Helly property
Our goal in this paper is to investigate the fractional Helly theorem for set systems that are more general than the set system of all convex subsets of $\mathbb{R}^{d}$. In such a setting it might not be clear what should play the role of the dimension, and we can not really hope to determine the exact dependency between $\alpha$ and $\beta$. Instead, we aim for existence results, which motivates the following definitions (initially introduced in [1]).
We say that a set system $\cal C$ satisfies the fractional Helly property if there exist an integer $k$ and a function $\beta:(0,1)\to(0,1)$ for which the following holds: Every finite family $F\subset{\cal C}$ which contains at least $\alpha\cdot\binom{|F|}{k}$ intersecting $k$-tuples, also contains an intersecting subfamily of size at least $\beta(\alpha)\cdot|F|$. In this situation we also say that $\cal C$ satisfies the fractional Helly property for $k$-tuples. The smallest integer $k$ such that $\cal C$ satisfies the fractional Helly property for $k$-tuples (if it exists) is called the fractional Helly number for $\cal C$.
It is important to note that the fractional Helly property is trivially satisfied if $\cal C$ is finite, and so it is really only interesting to consider the fractional Helly property for infinite set systems. A concrete example of an infinite set system is when $\cal C$ is the set system of all convex sets in $\mathbb{R}^{d}$, in which case the fractional Helly number equals $d+1$.
1.3. Convex lattice sets
Let us review another, rather remarkable, example. A subset $S\subset\mathbb{Z}^{d}$ is a convex lattice set if there exists a convex set $C\subset\mathbb{R}^{d}$ such that $S=C\cap\mathbb{Z}^{d}$. There is a Helly theorem for finite families of convex lattice sets, due to Doignon [7], but it requires that every $2^{d}$ or fewer members of the family have non-empty intersection. We note that the theorem is sharp in the sense that $2^{d}$ cannot be replaced by a strictly smaller value.
Regarding the fractional Helly property, it was observed in [1, section 8] that the method of Katchalski and Liu (a collapsibility argument) can be used to show that the set system of all convex lattice sets in $\mathbb{Z}^{d}$ satisfies the fractional Helly property for $2^{d}$-tuples. However, the following result, due to Bárány and Matoušek [6], shows that the fractional Helly number is actually much smaller.
Theorem 1 (Bárány–Matoušek).
The fractional Helly number for the set system of all convex lattice sets in $\mathbb{Z}^{d}$ equals $d+1$.
Bárány and Matoušek’s result shows that the number $2^{d}$ is in a sense a local anomaly (necessary for a Helly theorem), while for more global properties the relevant number remains $d+1$ (sufficient for a fractional Helly theorem). Bárány and Matoušek’s proof relies mostly on tools from extremal combinatorics, and towards the end of their paper [6, page 234] they make the following remark:
“Our proof of the fractional Helly theorem for convex lattice sets does not use much of the geometric properties of $\mathbb{Z}^{d}$ It would be interesting to clarify what axioms are sufficient … say in the context of abstract convexity spaces."
The purpose of this paper is to address the remark of Bárány and Matoušek, and thereby extend their theorem to the setting of abstract convexity spaces.
1.4. Convexity spaces
A convexity space $(X,\mathcal{C})$ consists of a non-empty set $X$ and a collection of subsets $\mathcal{C}\subset 2^{X}$ satisfying the following properties.
  • (C1)
    $\emptyset$ and $X$ are in $\mathcal{C}$.
  • (C2)
    $\mathcal{C}$ is closed under intersections, that is, if $\mathcal{D}\subset\mathcal{C}$ is non-empty, then $\bigcap_{C\in\mathcal{D}}C$ is in $\mathcal{C}$.
  • (C3)
    $\mathcal{C}$ is closed under nested unions, that is, if $\mathcal{D}\subset\mathcal{C}$ is non-empty and totally ordered by inclusion, then $\bigcup_{C\in\mathcal{D}}C$ is in $\mathcal{C}$.
In a convexity space $(X,\mathcal{C})$ the members of $\mathcal{C}$ are called convex sets. Given a subset $Y\subset X$ we define the convex hull of $Y$, denoted by $\mbox{conv}(Y)$, as the intersection of all convex sets that contain $Y$.
1.5. Radon and Helly numbers
There are several parameters which measure the complexity of a convexity space. For us the most important parameter is the Radon number, which is defined to be the smallest integer $n$ (if it exists) such that any subset $Y\subset X$, with $|Y|=n$, admits a partition $Y=A\cup B$ such that $\mbox{conv}(A)\cap\mbox{conv}(B)\neq\emptyset$.
A closely related parameter is the Helly number. This is the smallest integer $n$ (if it exists) such that every finite family $F\subset\mathcal{C}$, with $\bigcap_{S\in F}S=\emptyset$, contains a subfamily $G\subset F$, with $|G|\leq n$ and $\bigcap_{S\in G}S=\emptyset$. The relationship between the Radon and Helly numbers of a convexity space is given by the following fundamental theorem of Levi [15].
Theorem 2 (Levi).
For any convexity space $(X,\mathcal{C})$ with bounded Radon number, the Helly number exists and is strictly less than the Radon number.
A more recent result, due to Holmsen and Lee [12], relates the fractional Helly number to the Radon number of a convexity space.
Theorem 3 (Holmsen–Lee).
For any convexity space $(X,\cal C)$ with bounded Radon number, the set system $\cal C$ has the fractional Helly property. In particular, the fractional Helly number is bounded by some function depending only on the Radon number.
The bound on the fractional Helly number obtained in [12] is roughly $r^{r^{\lceil\log_{2}r\rceil}}$ where $r$ is the Radon number. Our main goal is to show that the bound on the fractional Helly number can be significantly reduced for convexity spaces that satisfy an additional separation axiom. In fact, in Corollary 6 we show that in separable convexity spaces the fractional Helly number is at most $2^{r}$.
1.6. Separability
Given a convexity space $(X,\mathcal{C})$, a halfspace [^1] Also known as hemispace. is a convex set $\gamma\in\mathcal{C}$ such that the complement $\bar{\gamma}=X\setminus\gamma\in\mathcal{C}$. Note that $\emptyset$ and $X$ are both considered to be halfspaces. The convexity space $(X,\mathcal{C})$ is called separable if for any $S\in\mathcal{C}$ and $x\in X\setminus S$ there exists a halfspace $\gamma\in\mathcal{C}$ such that $S\subset\gamma$ and $x\in\bar{\gamma}$.222 In general there exists a hierarchy of “separation axioms”, and what we define here as separable is referred to as $S_{3}$-separability in [21, Chapter I.3].
Our goal is to show that for a separable convexity space with bounded Radon number, the fractional Helly number can be bounded by a suitable complexity measure on the set system of its halfspaces, which is typically much better than the bound guaranteed by Theorem 3.
1.7. Dual VC-dimension
Let $\cal F$ be a set system on a ground set $X$. We denote by $A$ the incidence matrix of $X$ and $\cal F$, where the rows are indexed by elements of $\cal F$ and columns by elements of $X$. The dual shatter function of $\cal F$ is a function $\pi_{\cal F}^{*}:\mathbb{N}\to\mathbb{N}$, where $\pi_{\cal F}^{*}(m)$ is the maximum number of distinct columns in $A^{\prime}$, where $A^{\prime}$ is an $m$-row submatrix of $A$ (the maximum is taken over all $m$-row submatrices). Note that $\pi_{\cal F}^{*}(m)$ counts the maximum number of nonempty regions of the Venn diagram of $m$ sets of $\cal F$.
The dual VC-dimension of $\cal F$ is the maximum possible number of sets in $\cal F$ with a complete Venn diagram, that is, $\max\{k:\pi_{\cal F}^{*}(k)=2^{k}\}$. The following result, due to Matoušek [18], establishes a connection between the dual VC-dimension of a set system $\cal F$ and the fractional Helly property.
Theorem 4 (Matoušek).
Let $\cal F$ be a set system with dual VC dimension $d$. Then $\cal F$ has fractional Helly number at most $d+1$.
1.8. Main result
We are now in position to state our main result.
Theorem 5.
Let $(X,{\cal C})$ be a separable convexity space with bounded Radon number, and suppose the system of halfspaces has dual VC-dimension $d$. Then the fractional Helly number for $\cal C$ is at most $d+1$.
It is instructive to compare this theorem to Matoušek’s result (Theorem 4), which gives a bound on the fractional Helly number for any set system $\cal F$ with bounded dual VC-dimension. Our result, however, can be viewed as a statement about symmetric set systems $\cal F$ ($\gamma\in\cal F\Rightarrow\bar{\gamma}\in\cal F$). While we obtain the same bound on the fractional Helly number in terms of the dual VC-dimension, the main strength of our result is that the fractional Helly property applies not only to $\cal F$, but to the entire set system =$\cal F^{\cap}=\big{\{}\bigcap_{A\in\cal G}A:\cal G\subset\cal F\big{\}}$. For instance, the set system of convex sets in $\mathbb{R}^{d}$ does not have bounded (dual) VC-dimension, thus Theorem 4 does not imply the fractional Helly property for convex sets in $\mathbb{R}^{d}$. On the other hand, set system of convex sets in $\mathbb{R}^{d}$ is separable [16] and it is straightforward to check that the dual VC-dimension of the system of halfspaces in $\mathbb{R}^{d}$ equals $d$. Indeed, first note that the number of regions in the complement of $m$ hyperplanes in $\mathbb{R}^{d}$ is at most $\sum_{i=0}^{d}\binom{m}{i}$ and equality is attained for hyperplanes in general position. It follows from [16] that closures of halfspaces [^3] Note that halfspace is in the abstract convexity setting more general then a standard affine halfspace. e.g. a maximal (with respect to inclusion) convex set in $\mathbb{R}^{d}$ not containing a single point $x$ is not an affine halfspace, but it is a halfspace in the abstract setting. in $\mathbb{R}^{d}$ are standard closed affine halfspaces, hence the regions in the complement of $m$ hyperplanes in $\mathbb{R}^{d}$ are in 1-1 correspondence with regions of the Venn diagram of the system of $m$ halfspaces in $\mathbb{R}^{d}$. A complete Venn diagram is obtained for $m\leq d$ hyperplanes in general position, hence, the dual VC-dimension of the system of halfspaces in $\mathbb{R}^{d}$ is indeed $d$. Theorem 5 implies that convex sets in $\mathbb{R}^{d}$ have fractional Helly number (at most) $d+1$.
It is easy to see that Bárány and Matoušek’s result on convex lattice sets (Theorem 1) is a special case of Theorem 5 as well. The set system of convex lattice sets in $\mathbb{Z}^{d}$ is separable where the halfspaces have the form $\mathbb{Z}^{d}\cap\gamma$ where $\gamma$ is a halfspace in $\mathbb{R}^{d}$. The dual VC-dimension of halfspaces in $\mathbb{Z}^{d}$ is $d$ (for the same reasons as in the case of $\mathbb{R}^{d}$), and a result due to Onn [20] shows that the Radon number for convex lattice sets in $\mathbb{Z}^{d}$ is at most $d(2^{d}-1)+3$, and so the result follows.
One might be tempted to ask if in a separable convexity space with bounded Radon number and in which the system of hyperplanes has dual VC-dimension $d$, the Helly number of convex sets is at most $d+1$. However, this is not true as the set system of convex lattice sets in $\mathbb{Z}^{d}$ satisfies the assumptions but has Helly number $2^{d}$ [7].
The complexity of the system of halfspaces of a separable convexity space was previously investigated by Moran and Yehudayoff [19], who used the VC-dimension (primal, not the dual) to establish a weak $\varepsilon$-net theorem. Let us recall that (primal) VC-dimension of a set system $\cal F$ is defined as $\max\{k:\pi_{\cal F}(k)=2^{k}\}$, where $\pi_{\cal F}(m)$ is defined as maximum number of distinct rows in any $m$-column submatrix of the incidence matrix $A$ of $X$ and $\mathcal{F}$ (recall that rows of $A$ are indexed by members of $\mathcal{F}$ and columns by elements of $X$) [^4] Compare with the definition of the dual shatter function. By [3] (for proof in English see [17, Lemma 10.3.4]),
[…]
Moran and Yehudayoff showed [19, Lemma 1.5] that in a separable convexity space with Radon number $r$, the family of halfspaces has VC-dimension at most $r-1$. Combining it with (1) and Theorem 5, we conclude the following:
Corollary 6.
Let $r\geq 3$ and $(X,\mathcal{C})$ be a separable convexity space with Radon number $r$. Then the fractional Helly number for $\mathcal{C}$ is at most $2^{r}$.
The bound $2^{r}$ in Corollary 6 is (asymptotically) tight. To see this, consider the example of box convexity on $\mathbb{R}^{d}$. This is the subspace of the standard convexity whose convex sets consists of axis-parallel boxes in $\mathbb{R}^{d}$, that is, Cartesian products of segments parallel to the coordinate axes. This space is separable; the halfspaces are bounded by axis parallel hyperplanes. The Radon number of this space equals
$\min\left\{r:\binom{r}{\lfloor r/2\rfloor}>2d\right\},$
which grows asymptotically as $\Theta(\log_{2}d)$. (See [9, section 3] for more details.) The fractional Helly number for box convexity in $\mathbb{R}^{d}$ is obviously bounded above by $d+1$, by the standard fractional Helly theorem, and it is easily seen that this can not be reduced. Take for instance $n$ axis parallel hyperplanes split into $d$ groups of size $n/d$, where the members of the $i$th group are orthogonal to the $i$th coordinate axis. This is a family of $n$ boxes which contains $(n/d)^{d}\geq\frac{1}{e^{d}}\binom{n}{d}$ intersecting $d$-tuples, but where the largest intersecting subfamily has size $d$.
Finally, let us see why a bounded Radon number is necessary in Theorem 5. Generalizing the notion of standard convexity in $\mathbb{R}^{d}$, as suggested in [5, Conjecture 2.9 and Problem 2.10], for positive integers $k$ and $d$, let ${\cal B}_{k}^{d}$ denote the sets of solutions in $\mathbb{R}^{d}$ of polynomial inequalities of degree at most $k$. Note that $\cal B_{k}^{d}$ is a separable convexity space where the halfspaces are sets of solutions in $\mathbb{R}^{d}$ of a single polynomial inequality. It follows by a simple lifting argument that the system of halfspaces has dual VC-dimension bounded by a constant depending on $k$ and $d$. Indeed, by linearization, a single polynomial inequality becomes a linear inequality in $\mathbb{R}^{\binom{d+k}{d}}$ and as the dual VC-dimension of standard euclidean halfspaces in $\binom{d+k}{d}$-dimensional space is $\binom{d+k}{d}$, the result follows.
However, for any numbers $a_{1}<a_{2}<\cdots<a_{n+1}$, the system of linear and quadratic inequalities
$\{a_{1}\leq x\leq a_{n+1}\}\cup\big{\{}0\leq y\leq(x-a_{i})(x-a_{i+1})\big{\}}_{i=1}^{n},$
has the points $(a_{1},0),(a_{2},0),\dots,(a_{n+1},0)$ as its set of solutions, and therefore $\cal B^{2}_{2}$ contains any finite subset of the $x$-axis. Consequently, for any family of finite sets there exists an isomorphic copy as a subfamily of $\cal B^{2}_{2}$, which implies that the Helly, Radon, and fractional Helly numbers are all unbounded. This disproves Conjecture 2.9 of Bárány and Kalai [5].
Furthermore, the previous example combined with (1) directly gives that there is no affirmative answer to Problem 2.10 in [5] unless one adds an additional assumption of bounded Radon number. Indeed, it follows from Theorem 5 and expression (1) that this additional assumption is sufficient: if the Radon number of a separable convexity space $(X,\mathcal{C})$ is bounded and the set of halfspaces has VC-dimension at most $d$, than the fractional Helly number for $\mathcal{C}$ is bounded by a function of $d$, which gives an affirmative answer to [5, Problem 2.10(i)].
2. Proof of the main result
Our proof of Theorem 5 relies on a “weak” colorful Helly theorem for separable convexity spaces with bounded Radon number. This is a generalization of a result of Bárány and Matoušek [6, Proposition 3.1] which deals with the case of convex lattice sets in $\mathbb{Z}^{d}$.
Proposition 7.
For any integers $r\geq 3$, $d\geq 1$, and $m\geq 2$ there exists a positive integer $p=p(r,d,m)$ with the following property. Let $(X,\mathcal{C})$ be a separable convexity space with Radon number at most $r$ and suppose the system of halfspaces has dual VC-dimension at most $d$. If we are given families $F_{1},\dots,F_{d+1}\subset\mathcal{C}$, with $|F_{i}|=p$, such that $\bigcap_{i=1}^{s}C_{i}\neq\emptyset$ for every choice $C_{1}\in F_{1},\dots,C_{d+1}\in F_{d+1}$, then one of the families $F_{i}$ contains $m$ members with non-empty intersection.
The proof closely follows the argument of Bárány and Matoušek. This is the main technical step of this paper and is given in Section 3.
With Proposition 7 at hand, Theorem 5 now follows from a standard counting argument using the Erdős–Simonovits supersaturation theorem [10]. Its use appears frequently in discrete geometry and a simple proof can be found in [17, Theorem 9.2.2]. For completeness we spell out the proof of Theorem 5 below.
For integers $k,s\geq 2$, let $K^{k}(s)$ denote the $k$-uniform $k$-partite hypergraph with each class of size $s$.
Proposition 8 (Erdős–Simonovits).
For all integers $k,s\geq 2$ and a real number $\varepsilon\in(0,1]$ there exists a $\delta>0$ with the following property. Let $H$ be a $k$-uniform hypergraph on $n$ vertices and at least $\varepsilon\binom{n}{k}$ edges. Then $H$ contains at least $\lfloor\delta n^{ks}\rfloor$ copies of $K^{k}(s)$.
Let us remark on an important special case: For any fixed $\varepsilon>0$, if a $k$-uniform hypergraph $H$ contains an $\varepsilon$-fraction of all possible edges and $n$ is sufficiently large in terms of $k$ and $\varepsilon$, then $H$ contains $K^{k}(s)$ as a subhypergraph (not necessarily induced).
Proof of Theorem 5.
Consider a separable convexity space $(X,\mathcal{C})$. Let $r$ be the Radon number and $d$ the dual VC-dimension of the system of halfspaces. Let $m$ be bound on the fractional Helly number given by Proposition 3 and let $p=p(r,d,m)$ be the integer from Proposition 7. Our goal is to show that any finite family $F\subset\cal C$ containing many intersecting $(d+1)$-tuples, will also contain many intersecting $m$-tuples.
Now suppose we are given a family $F\subset\mathcal{C}$, with $|F|=n$, which contains at least $\alpha\binom{n}{d+1}$ intersecting $(d+1)$-tuples. Let $H$ be a $(d+1)$-uniform hypergraph on $n$ vertices which are in one-to-one correspondence with the members of $F$, and let a $(d+1)$-tuple of vertices of $H$ form an edge whenever the corresponding $(d+1)$-tuple of $F$ is intersecting. Thus $H$ has at least $\alpha\binom{n}{d+1}$ edges, and so by the Erdős–Simonovits theorem (Poposition 8) $H$ contains at least $\delta n^{p(d+1)}$ distinct copies of $K^{d+1}(p)$ for some $\delta>0$ (which depends only on $\alpha$, $d$, and $p$). Note also that a fixed $p(d+1)$-tuple of vertices can contain at most a constant number (depending only on $p$ and $d$) of distinct copies of $K^{d+1}(p)$, and so $H$ contains at least $\delta^{\prime}n^{p(d+1)}$ distinct copies of $K^{d+1}(p)$ each supported on a distinct $p(d+1)$-tuple of vertices, for some $\delta^{\prime}>0$ (which still only depends on $\alpha$, $d$, and $p$). By Proposition 7, for each such copy we find an $m$-tuple of $F$ with non-empty intersection. Each such intersecting $m$-tuple will be counted in this way by at most $\binom{n-m}{p(d+1)-m}$ distinct copies of $K^{d+1}(p)$, which means there are at least
${\frac{\delta^{\prime}n^{p(d+1)}}{\binom{n-m}{p(d+1)-m}}\geq cn^{m}\geq\alpha^{\prime}\binom{n}{m}}$
distinct $m$-tuples of $F$ with non-empty intersection, where $\alpha^{\prime}>0$ depends only on $\alpha$, $d$, and $p$. We may now apply Proposition 3 to find $\beta n$ members of $F$ with non-empty intersection, where $\beta=\beta(\alpha^{\prime},r)>0$. ∎
3. Auxiliary results
As mentioned above, the proof of Proposition 7 is the main technical step of this paper. The proof relies on two auxiliary results regarding collections of point sets in a separable convexity space. These results (Lemmas 9 and 10 below) were originally proved by Bárány and Matoušek in the setting of convex lattice sets in $\mathbb{Z}^{d}$. Surprisingly, in our extensions, the dual VC-dimension of the system of halfspaces does not play any role, and we require only that the Radon number is bounded. The bound on the dual VC-dimension only comes into play in the very last step of the proof of Proposition 7.
Let $X$ be the ground set of a convexity space. Throughout this section we frequently deal with finite multisets of points from $X$. It will be convenient to think of a multiset as the image of a function $f:E\to X$, where $E$ is a non-empty finite set. The multiplicity of a point $x\in X$ is defined as the cardinality of its preimage, that is, $|f^{-1}(x)|$. For a finite subset $Z\subset X$, we write $|Z|_{f}=\sum_{x\in Z}|f^{-1}(x)|$. Here is an extension of [6, Corollary 2.2].
Lemma 9.
Let $(X,\mathcal{C})$ be a separable convexity space with Radon number at most $r$, let $E$ be a non-empty finite set, and let $m\geq 2$ be an integer. For any finite family of functions $\{f_{i}\}_{i=1}^{m}$, where $f_{i}:E\to X$ and $\bigcap_{i=1}^{m}\mbox{\em conv}(f_{i}(E))=\emptyset$, there exist
    a subset $E_{0}\subset E$, with $|E_{0}|\geq\frac{1}{r-1}|E|$,
    a pair of functions $f_{i}$ and $f_{j}$, and
    a halfspace $\gamma\in\mathcal{C}$,
such that the images $f_{i}(E_{0})\subset\gamma$ and $f_{j}(E_{0})\subset\bar{\gamma}$.
Proof.
Let $S=\bigcup_{i=1}^{m}f_{i}(E)$ and set $N=\sum_{i=1}^{m}|S|_{f_{i}}=m|E|$. Define the family
$F=\left\{\>{\mbox{conv}(Z)\;:\;Z\subseteq S},\>\textstyle{\sum_{i=1}^{m}|Z|_{f_{i}}>\frac{r-2}{r-1}}N\right\}.$
Note that $F$ is finite as $S$ is finite and nonempty since $\mbox{conv}(S)\in F$. Also note that every $k\leq r-1$ members of $F$ have a point in common as the cardinality of the preimages of their complements is strictly smaller than $N$ (we use that $Z\subseteq\mbox{conv}(Z))$. We claim that there exists a point $x_{0}$ common to all members of $F$. Indeed, if $|F|\leq r-1$ it is trivial, otherwise it follows from Theorem 2.
By assumption there is an $f_{i}$ such that $x_{0}\not\in\mbox{conv}(f_{i}(E))$, and so by separability there is a halfspace $\gamma$ in $\cal C$ such that
$f_{i}(E)\subseteq\mbox{conv}(f_{i}(E))\subseteq\gamma\>\mbox{ and }\>x_{0}\in\bar{\gamma}.$
Moreover, as $x_{0}$ belongs to every set in $F$ but it is not contained in the halfspace $\gamma$, we have $\gamma\notin F$, hence ${\sum_{j=1}^{m}|\gamma|_{f_{j}}\leq\frac{r-2}{r-1}N}$, which implies $\sum_{j=1}^{m}|\bar{\gamma}|_{f_{j}}\geq\frac{N}{r-1}$. Note also that $|\bar{\gamma}|_{f_{i}}=0$ since $f_{i}(E)\subseteq\gamma$. By the pigeon-hole principle there is a $j\in[m]\setminus\{i\}$ such that
${|\bar{\gamma}|_{f_{j}}\geq\frac{N}{(m-1)(r-1)}=\frac{m|E|}{(m-1)(r-1)}\geq\frac{1}{r-1}|E|}.$
To finish the proof, we set $E_{0}=\{e\in E\>:\>f_{j}(e)\in\bar{\gamma}\}$. ∎
Let $H=(V,E)$ be a $k$-uniform hypergraph. For a vertex $v\in V$ and a subset $\tilde{E}\subset E$, we define the subset $\tilde{E}_{v}\subset\tilde{E}$ as
$\tilde{E}_{v}=\{e\in\tilde{E}\>:\>v\in e\}.$
Suppose we are given a separable convexity space $(X,\mathcal{C})$, a function $f:E\to X$, and a subhypergraph $\tilde{H}=(\tilde{V},\tilde{E})\subset H$. A subset of vertices $A\subset\tilde{V}$ is called separable with respect to $\tilde{H}$ if there exists
    a pair of vertices $u,v\in A$, and
    a halfspace $\gamma\in\mathcal{C}$
with the property that
$f(\tilde{E}_{u})\subset\gamma\>\mbox{ and }\>f(\tilde{E}_{v})\subset\bar{\gamma}.$
We now have the following extension of [6, Lemma 3.2].
Lemma 10.
For any integers $n\geq m\geq 2$, $k\geq 2$ and $r\geq 3$ there exists an integer $t=t(n,m,k,r)$ with the following property. Let $(X,\mathcal{C})$ be a separable convexity space with Radon number at most $r$, and let $H=K^{k}(t)$ with vertex classes $V_{1},\dots,V_{k}$ and edge set $E$. For any function $f:E\to X$, one of the following hold:
  1. (1)
    There is an $m$-element set $A\subset V_{k}$ such that
    ${\bigcap_{v\in A}}\mbox{\em conv}(f(E_{v}))\neq\emptyset.$
  2. (2)
    There are $n$-element sets $W_{1}\subset V_{1},\dots,W_{k}\subset V_{k}$ such that every $m$-element subset of $W_{k}$ is separable with respect to the induced subhypergraph $H[\bigcup_{i=1}^{k}W_{i}]$.
Proof.
Fix an arbitrary $n$-element set $W_{k}\subset V_{k}$, and consider an $m$-element subset $A\subset W_{k}$. Suppose that $A$ does not fall into case (1), that is,
${\bigcap_{v\in A}}\mbox{conv}(f(E_{v}))=\emptyset.$
We claim that for any integer $s$, if $t$ is sufficiently large, then there will exists $s$-element subsets $U_{1}\subset V_{1},\dots,U_{k-1}\subset V_{k-1}$ such that $A$ is separable with respect to the induced subhypergraph $H[U_{1}\cup\cdots\cup U_{k-1}\cup W_{k}]$.
Note that this claim will prove the lemma, because after applying the claim at most $\binom{n}{m}$ times and passing to smaller $U_{i}$’s at each step, we can make every $m$-element subset of $W_{k}$ separable with respect to the resulting induced subhypergraph $H[\bigcup_{i=1}^{k}W_{i}]$. We can guarantee that $|W_{1}|=\cdots=|W_{k-1}|=n$ by starting with $t$ sufficiently large.
In order to prove the claim we are going to apply Lemma 9. Let $\tilde{E}$ denote the edge set of the $(k-1)$-uniform hypergraph $K^{k-1}(t)$ on the vertex classes $V_{1},\dots,V_{k-1}$. For every $v\in A$, define a function $f_{v}:\tilde{E}\to X$ by setting $f_{v}(\sigma)=f(\sigma\cup\{v\})$. Note that $f_{v}(\tilde{E})=f(E_{v})$. By Lemma 9 there exist
    a subset $\tilde{E}_{0}\subset\tilde{E}$, with $|\tilde{E}_{0}|\geq\frac{1}{r-1}|\tilde{E}|$,
    a pair of vertices $u,v\in A$, and
    a halfspace $\gamma\in\mathcal{C}$,
such that $f_{u}(\tilde{E}_{0})\subset\gamma$ and $f_{v}(\tilde{E}_{0})\subset\bar{\gamma}$.
To complete the proof of the claim, we apply the Erdős–Simonovits theorem. Since $\tilde{E}_{0}$ contains at least a $\frac{1}{r-1}$-fraction of the edges of $K^{k-1}(t)$ and hence at least an $\varepsilon$-fraction of all possible edges of a $(k-1)$-uniform hypergraph on $t(k-1)$ vertices, where $\varepsilon$ depends on $k$ and $r$, it follows that if we choose $t$ sufficiently large, then $\tilde{E}_{0}$ will contain the edge set of a copy of $K^{k-1}(s)$. Setting $U_{1},\dots,U_{k-1}$ to be the vertex classes of this $K^{k-1}(s)$ proves the claim. ∎
4. Proof of Proposition 7
Recall that $F_{1},\dots,F_{d+1}$ are families of convex sets in a separable convexity space $(X,\mathcal{C})$ with Radon number at most $r$ and where the system of halfspaces has dual VC-dimension at most $d$. For every $i\in[d+1]$ we have $|F_{i}|=p$, and for every choice $C_{1}\in F_{1},\dots,C_{d+1}\in F_{d+1}$ we have $C_{1}\cap\cdots\cap C_{d+1}\neq\emptyset$. For contradiction, let us assume that for every $i\in[d+1]$, each $m$-membered subfamily of $F_{i}$ has empty intersection.
Let $H=K^{d+1}(p)$ with vertex classes $V_{1},\dots,V_{d+1}$ and edge set $E$. For every vertex $v\in V_{i}$ we associate a convex set $C(v)\in F_{i}$, so that for every $i\in[d+1]$ we have a bijection between the vertices in $V_{i}$ and the members of $F_{i}$. Define a function $f:E\to X$ by (arbitrarily) choosing a point
${f(e)\in\bigcap_{v\in e}C(v)},$
which is possible by our hypothesis.
For sufficiently large $p$, we may apply Lemma 10 to obtain an induced subhypergraph such that every $m$-tuple of the $(d+1)$st vertex class is separable with respect to this subhypergraph. Assuming $p$ was even larger when we started, this can be repeated another $d$ times (relabeling the vertex classes each time), eventually ending up with an induced subhypergraph $H^{\prime}=K^{d+1}(2)$ with vertex classes $W_{1},\dots,W_{d+1}$ and edge set $\tilde{E}$ such that each vertex class is separable with respect to $H^{\prime}$. In particular, if we let $W_{i}=\{u_{i},v_{i}\}$, then there exists a halfspace $\gamma_{i}\in\mathcal{C}$ such that $f(\tilde{E}_{u_{i}})\subset\gamma_{i}$ and $f(\tilde{E}_{v_{i}})\subset\bar{\gamma}_{i}$. But this implies that for $e=\{a_{1},\dots,a_{d+1}\}\in\tilde{E}$ we have
$f(e)\in A_{1}\cap\cdots\cap A_{d+1},$
where
$A_{i}=\begin{cases}\gamma_{i}&\mbox{when }a_{i}=u_{i}\\ \bar{\gamma}_{i}&\mbox{when }a_{i}=v_{i}.\end{cases}$
But this implies that the halfspaces $\gamma_{1},\dots,\gamma_{d+1}$ have a complete Venn diagram, which contradicts the assumption on the dual VC-dimension of the system of halfspaces. ∎
5. Acknowledgements
The work was partially supported by ERC Advanced Grants “GeoScape” 882971 and “ERMiD” 101054936. A. H. was supported by the Institute for Basic Science (IBS-R029-C1), Z. P. was supported by the GAČR grant no. 22-19073S. Part of the work was done during the Erdős Center Focused week: Combinatorial geometry in Radon convexity spaces, which took place in Budapest in December 2023.

### v3

@@
 a maximal (with respect to inclusion) convex set in $\mathbb{R}^{d}$ not containing a single point $x$ is not an affine halfspace, but it is a halfspace in the abstract setting.
-in $\mathbb{R}^{d}$ are standard closed affine halfspaces, hence the regions in the complement of $m$ hyperplanes in $\mathbb{R}^{d}$ are in 1-1 correspondence with regions of the Venn diagram of the system of $m$ halfspaces in $\mathbb{R}^{d}$.
+in $\mathbb{R}^{d}$ are standard closed affine halfspaces, hence the regions in the complement of $m$ hyperplanes in $\mathbb{R}^{d}$ are in one-to-one correspondence with regions of the Venn diagram of the system of $m$ halfspaces in $\mathbb{R}^{d}$.
 A complete Venn diagram is obtained for $m\leq d$ hyperplanes in general position, hence, the dual VC-dimension of the system of halfspaces in $\mathbb{R}^{d}$ is indeed $d$.
@@
 Finally, let us see why a bounded Radon number is necessary in Theorem 5.
-Generalizing the notion of standard convexity in $\mathbb{R}^{d}$, as suggested in [5, Conjecture 2.9 and Problem 2.10], for positive integers $k$ and $d$, let ${\cal B}_{k}^{d}$ denote the sets of solutions in $\mathbb{R}^{d}$ of polynomial inequalities of degree at most $k$.
+Generalizing the notion of standard convexity in $\mathbb{R}^{d}$, as suggested in [5, Conjecture 2.9 and Problem 2.10], for positive integers $k$ and $d$, let ${\cal B}_{k}^{d}$ denote the sets of solutions in $\mathbb{R}^{d}$ of polynomial inequalities (strict and nonstrict) of degree at most $k$.
 Note that $\cal B_{k}^{d}$ is a separable convexity space where the halfspaces are sets of solutions in $\mathbb{R}^{d}$ of a single polynomial inequality.
 It follows by a simple lifting argument that the system of halfspaces has dual VC-dimension bounded by a constant depending on $k$ and $d$.
-Indeed, by linearization, a single polynomial inequality becomes a linear inequality in $\mathbb{R}^{\binom{d+k}{d}}$ and as the dual VC-dimension of standard euclidean halfspaces in $\binom{d+k}{d}$-dimensional space is $\binom{d+k}{d}$, the result follows.
-However, for any numbers $a_{1}<a_{2}<\cdots<a_{n+1}$, the system of linear and quadratic inequalities
+Indeed, by linearization, a single polynomial inequality becomes a linear inequality in $\mathbb{R}^{\binom{d+k}{d}}$ and as the dual VC-dimension of standard euclidean halfspaces in $\binom{d+k}{d}$-dimensional space equals $\binom{d+k}{d}$, the claim follows.
+On the other hand, the convexity space ${\cal B}^{d}_{k}$ is in a sense “universal” already for $d=k=2$.
+Proposition 7.
+Let =$\cal F=\{S_{1},S_{2},\dots,S_{m}\}$ be a set system on a ground set $X$.
+There exists a family of convex sets $\{C_{1},C_{2},\dots,C_{m}\}\subset{\cal B}^{2}_{2}$ such that for every $I\subset[m]$ we have $\bigcap_{i\in I}S_{i}\neq\emptyset$ if and only if $\bigcap_{i\in I}C_{i}\neq\emptyset$.
+Proof.
+For any numbers $a_{1}<a_{2}<\cdots<a_{n+1}$, the system of linear and quadratic inequalities
 $\{a_{1}\leq x\leq a_{n+1}\}\cup\big{\{}0\leq y\leq(x-a_{i})(x-a_{i+1})\big{\}}_{i=1}^{n},$
 has the points $(a_{1},0),(a_{2},0),\dots,(a_{n+1},0)$ as its set of solutions, and therefore $\cal B^{2}_{2}$ contains any finite subset of the $x$-axis.
-Consequently, for any family of finite sets there exists an isomorphic copy as a subfamily of $\cal B^{2}_{2}$, which implies that the Helly, Radon, and fractional Helly numbers are all unbounded.
+More precisely, we index non-empty regions of the Venn diagram $V(\cal F)$ of $\cal F$ by distinct numbers $a_{j}$, and for every $i\in[m]$, we let $C_{i}$ consist of all $(a_{j},0)$ where $a_{j}$ is assigned to a region of $V(\cal F)$ contained in $S_{i}$.
+Clearly, intersection patterns of $\cal F$ are in one-to-one correspondence with intersection patterns of $\{C_{1},\ldots,C_{m}\}$.
+∎
+An immediate consequence of Proposition 7 is that the Helly, Radon, and fractional Helly numbers of ${\cal B}_{2}^{2}$ are all unbounded.
+[^5] Since there exist set systems for which Radon, Helly, and fractional Helly numbers, respectively, are unbounded.
 This disproves Conjecture 2.9 of Bárány and Kalai [5].
-Furthermore, the previous example combined with (1) directly gives that there is no affirmative answer to Problem 2.10 in [5] unless one adds an additional assumption of bounded Radon number.
+Furthermore, Proposition 7 combined with (1) directly gives that there is no affirmative answer to Problem 2.10 in [5] unless one adds an additional assumption of bounded Radon number.
 Indeed, it follows from Theorem 5 and expression (1) that this additional assumption is sufficient: if the Radon number of a separable convexity space $(X,\mathcal{C})$ is bounded and the set of halfspaces has VC-dimension at most $d$, than the fractional Helly number for $\mathcal{C}$ is bounded by a function of $d$, which gives an affirmative answer to [5, Problem 2.10(i)].
@@
 Let $r$ be the Radon number and $d$ the dual VC-dimension of the system of halfspaces.
-Let $m$ be bound on the fractional Helly number given by Proposition 3 and let $p=p(r,d,m)$ be the integer from Proposition 7.
+Let $m$ be bound on the fractional Helly number given by Theorem 3 and let $p=p(r,d,m)$ be the integer from Proposition 8.
 Our goal is to show that any finite family $F\subset\cal C$ containing many intersecting $(d+1)$-tuples, will also contain many intersecting $m$-tuples.
@@
 Let $H$ be a $(d+1)$-uniform hypergraph on $n$ vertices which are in one-to-one correspondence with the members of $F$, and let a $(d+1)$-tuple of vertices of $H$ form an edge whenever the corresponding $(d+1)$-tuple of $F$ is intersecting.
-Thus $H$ has at least $\alpha\binom{n}{d+1}$ edges, and so by the Erdős–Simonovits theorem (Poposition 8) $H$ contains at least $\delta n^{p(d+1)}$ distinct copies of $K^{d+1}(p)$ for some $\delta>0$ (which depends only on $\alpha$, $d$, and $p$).
+Thus $H$ has at least $\alpha\binom{n}{d+1}$ edges, and so by the Erdős–Simonovits theorem (Poposition 9) $H$ contains at least $\delta n^{p(d+1)}$ distinct copies of $K^{d+1}(p)$ for some $\delta>0$ (which depends only on $\alpha$, $d$, and $p$).
 Note also that a fixed $p(d+1)$-tuple of vertices can contain at most a constant number (depending only on $p$ and $d$) of distinct copies of $K^{d+1}(p)$, and so $H$ contains at least $\delta^{\prime}n^{p(d+1)}$ distinct copies of $K^{d+1}(p)$ each supported on a distinct $p(d+1)$-tuple of vertices, for some $\delta^{\prime}>0$ (which still only depends on $\alpha$, $d$, and $p$).
@@
 distinct $m$-tuples of $F$ with non-empty intersection, where $\alpha^{\prime}>0$ depends only on $\alpha$, $d$, and $p$.
-We may now apply Proposition 3 to find $\beta n$ members of $F$ with non-empty intersection, where $\beta=\beta(\alpha^{\prime},r)>0$.
+We may now apply Theorem 3 to find $\beta n$ members of $F$ with non-empty intersection, where $\beta=\beta(\alpha^{\prime},r)>0$.
 ∎
@@
 The proof relies on two auxiliary results regarding collections of point sets in a separable convexity space.
-These results (Lemmas 9 and 10 below) were originally proved by Bárány and Matoušek in the setting of convex lattice sets in $\mathbb{Z}^{d}$.
+These results (Lemmas 10 and 11 below) were originally proved by Bárány and Matoušek in the setting of convex lattice sets in $\mathbb{Z}^{d}$.
 Surprisingly, in our extensions, the dual VC-dimension of the system of halfspaces does not play any role, and we require only that the Radon number is bounded.